OpenStudy (anonymous):

How to do (x^2 - x - 1)(2x^2 + 3x + 4)?

OpenStudy (anonymous):

First?: \[x^2(2x^2)? = 2x^4?\]

OpenStudy (anonymous):

x^2*2x^2+x^2*3x+x^2*4 -x*2x^2 -x*3x -x*4 -1(2x^2+3x+4)

OpenStudy (anonymous):

yes x^2*x^2 = x^4

OpenStudy (anonymous):

Okay, and then: \[[1]x(3x)? = 3x\]

OpenStudy (anonymous):

no, x(3x) = 3x^2

OpenStudy (anonymous):

And: 1(4) = 4 So..\[2x^4+3x^2+4?\]

OpenStudy (anonymous):

yes if you add up the previous various you'll get that